Mr. D. Bum has 5 quarts of punch that is 53% lemonade. He wants to make some punch containing 82% lemonade. Mr. Bum can add either water or lemonade to the punch, but not both. How much water or lemonade should he add? L represents the amount of lemonade added. Type in an equation that solves for L.
Answer by mananth(16946)

percent ---------------- quantity
53 ---------------- 5
100 ---------------- L
82 ---------------- 5+L
53*5+100*L=82( 5+L)

265 + 100 L = 410 + 82 L

100 L - 82 L = 410 - 265

18 L = 145

/ 18

L = 8.06 quarts